OpenStudy (anonymous):

Which problem illustrates the identity property? 5+0=5 5x0=0

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

i think its "5x0=0' just wanna make sure tho

OpenStudy (skullpatrol):

You just asked and got an answer to this :-)

OpenStudy (anonymous):

so im correct?

OpenStudy (anonymous):

5 + 0 = 5

OpenStudy (anonymous):

its 5 + 0 = 5

OpenStudy (skullpatrol):

yes

OpenStudy (anonymous):

because 5*0 !=5

OpenStudy (anonymous):

imma take the majority vote on this one skullpatrol, thanks anyways, and thanks to you both!

OpenStudy (skullpatrol):

$$\Huge 0*a=0$$ is the multiplicative property of 0